John is 3 years older than Mary. Five years ago he was four times as old as Mary was then. How old is John?
Found 2 solutions by josgarithmetic, MathTherapy:
Answer by josgarithmetic(39617)
Answer by MathTherapy(10552)


John is 3 years older than Mary. Five years ago he was four times as old as Mary was then. How old is John?
Let John's age be J
Then Mary's is: J - 3
We then get: J - 5 = 4(J - 3 - 5)
J - 5 = 4(J - 8)
Continue to solve for J, John's current age.